OpenStudy (anonymous):

PLEASE PLEASE HELP ME I WILL FAN AND MEDAL What is the value of u + v + w when u = –10, v = 4, and w = 8? A. –22 B. 2 C. 14 D. 22

OpenStudy (cjn1130):

2 B

OpenStudy (cjn1130):

-10 + 4 = -6 -6 + 8 = 2

OpenStudy (anonymous):

are you sure? i thought you added?

OpenStudy (cjn1130):

yeah but you said u=-10, so you either subtract it from 12 (8+4) or start with it

OpenStudy (anonymous):

can i ask a few more?

OpenStudy (cjn1130):

yeah sure

OpenStudy (anonymous):

What is the value of the variable that makes the statement true? c + (–8) + (–5) = 0 A. –13 B. –3 C. 3 D. 13

OpenStudy (cjn1130):